At 0.0°C and 1.00 atm, 1.00 mol of a gas occupies a volume of 0.0224 m3.

Estimate the average distance between the molecules.


1
Expert's answer
2022-05-18T15:28:21-0400

"pV=\frac mM RT,\implies"

"m=\frac{MpV}{RT}=14.01~g."

"\rho=\frac mV=0.625~\frac{kg}{m^3}."